OpenStudy (anonymous):

Please help Express \(36x^2 - 252\) as a difference of two squares the answer is supposed to have some binomials..idk

Parth (parthkohli):

\( \color{Black}{\Rightarrow (6x + \sqrt{252})(6x - \sqrt{252}) }\) Simplify sqrt252 further and you're done. Can you do it now?

OpenStudy (anonymous):

252 is not a perfect square sadly ;) sorry

OpenStudy (anonymous):

Hint: the squares are binomials

Parth (parthkohli):

I know but do you know how to simplify radicals?

OpenStudy (anonymous):

252 = 15.87

Parth (parthkohli):

and it's your last day so live it because we'll have a verification system

OpenStudy (anonymous):

if you don't know the answer get out of the post ;)

OpenStudy (anas.p):

\[x=\sqrt{7}\] i think so.

OpenStudy (anonymous):

no i dont think so..the answer has some binomials

OpenStudy (anonymous):

@ParthKohli What do you mean by "it's your last day so live it because we'll have a verification system"? Just curious.

OpenStudy (anonymous):

it's a threat lol

OpenStudy (anonymous):

Alright, did you get your answer for this question yet though?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Well, do you remember how to simplify radicals such as √252?

OpenStudy (anonymous):

\[\sqrt{252} = \sqrt{9 * 4 * 7} = 3 * 2\sqrt{7} = 6\sqrt{7}\]? Does this look familiar?

OpenStudy (anonymous):

Well, you just plug that into what ParthKohli did. (6x + 6√7)(6x - 6√7)

OpenStudy (anonymous):

@mr.awesome Do yuo get it?

OpenStudy (anonymous):

actually the answer is the diffence of two squares of binomials..that's what i dont get

OpenStudy (anonymous):

ParthKohli has the correct answer mr.awesome, it is the difference of two squares of binomials...i noticed that many students are misled when teachers in the classroom make it seem as though you can only have difference of two squares when the TERM A& Term B of (TERM A) ^2 -(TERM B)^2 are whole numbers and perfect squares i.e. x^2-1, B=1 is a whole number and a perfect square and x^2-1=(x+1)(x-1) but then later on you run into situations when term A& B aren't whole numbers and you still have (TERM A) ^2 - (TERM B)^2= (TERM A + TERMB)(TERMA-TERMB) take for instance 37x^2-11= ( sqrt(37)x+ sqrt(11)) (sqrt(37)x-sqrt(11))